OpenStudy (anonymous):

Factor completely x2 − 8x + 16. (x + 4)(x + 4) (x − 4)(x − 4) (x + 4)(x − 4) (x − 2)(x − 8)

OpenStudy (er.mohd.amir):

x2-4x-4x+16 =x(x-4)-4(x-4) =(x-4)(x-4)

OpenStudy (anonymous):

THANK YA